how do I solve 8x - 4 = 7x
Answer by ikdeep(189)

8x - 4 = 7x
subtract 7x from both sides and we get....
8x - 7x - 4 = 0
x - 4 = 0
subtract 4 from both sides and we get ..........
x = 4
Now if you want to verify you answer, you can put the values of x in the given equation and if you get LHS = RHS, this means that you answer is correct..
Hope this will help you.
Please feel free to revert back for any further queries.